Question 9 Background :

Model for IQ Scores

The distribution of IQ scores (for a particular age group) is approximately bell-shaped with a mean of 50 points and a standard deviation of 15 points.   

Question:

One individual scored 68 points on this test. Which of the following is a correct statement about this individual's score?

  • The individual scored 18 standard deviations above the mean score.

  • The individual scored 18 standard deviations below the mean score.

  • The individual scored 1.2 standard deviations above the mean score.

  • The individual scored 1.2 standard deviations below the mean score.
Answer #1

We calculate z score corresponding to score 68

as z= (68-50)/15=18/15=1.2

Then correct interpretation of the above z score is

The individual scored 1.2 standard deviations above the mean score. (THIRD option is correct)